# Undefinability of Approximation of 2-to-2 Games

### v1

@@ 3.3 The final (weighted) 2-to-2 game
The final step of the reduction is to transform the transitive game constructed in Section 3.2 into a weighted $2$-to-$2$ game, getting rid of the $1$-to-$1$ constraints. This weighted game is defined as follows.
Recall the transitive $2$-to-$2$ game $\Theta(I)$ constructed in Section 3.2. The transitivity condition guarantees that the vertices of $\Theta(I)$ can be partitioned into cliques $C_{1},\dots,C_{m}$ so that edges in each clique are associated with $1$-to-$1$ constraints. Moreover, these constraints are consistent in the sense that any colouring of a vertex $V$ in a clique $C$ can be extended in a unique way to a colouring of all vertices in $C$ so that all edge constraints in $C$ are satisfied. Also, by the transitivity condition, for distinct cliques $C_{i}$ and $C_{j}$, either all pairs $(u,v)\in C_{i}\times C_{j}$ are connected by $2$-to-$2$ constraints or none are. Furthermore, these $2$-to-$2$ constraints are consistent in the sense that given a clique-consistent colouring for $C_{i}$ and $C_{j}$, either all or none of these $2$-to-$2$ constraints are satisfied.
@@
  • Choose a random pair $L,L^{\prime}$ so that $(U,L)$ and $(U,L^{\prime})$ are connected by a $2$-to-$2$ edge. Let $C_{i}$ be the clique containing $(U,L)$ and $C_{j}$ be the clique containing $(U^{\prime},L^{\prime})$
@@ 3.4 Irregular soundness case
For the result in Section 5.3, we need the $\mathrm{FPC}$-undefinability of a different gap problem based on $2$-to-$2$ games. Specifically, we define the value of a game to be, not the fraction of constraints that can be satisfied, but the fraction of the vertices formed by the largest set $X$ so that all constraints between nodes in $X$ are satisfied. Moreover, we relax the notion of colouring to allow vertices to be coloured by multiple colours.
@@
For a $2$-to-$2$ game $((V,E),\Sigma,\Phi)$, a colouring $c:V\rightarrow\binom{\Sigma}{j}$ satisfies a set $X\subseteq V$ if $\forall e\in E$, such that the vertices of $e$ are in $X$, $\exists a\in c(u),b\in c(v).(a,b)\in\Phi(e)$.
@@
[…]
We can now state the theorem below, which is a consequence of Theorem 3.1.
@@
For every $\delta$ with $10<\delta<$ 1 and $j\in\mathbb{N}^{+}$, there exists $q\in\mathbb{N}^{+}$ so that $\mathrm{Gap}\textsc{Irreg2to2}_{j,q}(1,\delta)$ is not $\mathrm{FPC}$ definable.
[…]
@@
[…]
@@
Let $c$ be a $j$-colouring of $V$ that satisfies a set $X$ with $|X|/|V|\geq\delta$. By [24, Remark 3.4.9], there is a $\Omega(\delta^{2})$ (weighted) fraction of the edges $E$ which are satisfied by $c$, in the sense that for each such edge $e$ on vertices $u,v$ there are colours $a$ and $b$ in $c(u)$ and $c(v)$ respectively such that $(a,b)\in\Phi(e)$. We now construct a standard colouring by a random process. That is, for each vertex $v\in V$, independently choose a colour $\chi(v)$ from $c(v)$ uniformly at random. For an edge $e$ on vertices $u,v$, let $\Xi(e)$ be the indicator variable indicating whether $(\chi(u),\chi(v))\in\Phi(e)$ and let $\Xi$ be the overall value of the colouring $\chi$. If $(u,v)\in X^{2}$, the probability that $\chi$ satisfies the constraint $\Phi(e)$ is at least $\frac{1}{j^{2}}$, as by definition, among the $j^{2}$ pairs in $c(u)\times c(v)$, at least one satisfies the constraint. Then, we have,
@@
$\geq\sum_{e\in E|_{X}}w(e)\mathbb{E}[\Xi(e)]\geq\sum_{e\in E|_{X}}w(e)\frac{1}{j^{2}}\geq\Omega(\delta^{2})\frac{1}{j^{2}}$
where $E|_{X}$ denotes the subset of $E$ containing those edges both of whose vertices are in $X$. Thus, there is a colouring that satisfies at least $\Omega(\frac{\delta^{2}}{j^{2}})$ (weighted) fraction of the constraints. ∎
[…]
@@ 3.5 $2\leftrightarrow 2$ games
The definition of $2$-to-$2$ games, Definition 2.4 only requires each constraint $\Phi(u,.v)$ to be a $2$-to-$2$ relation, meaning that each element on the left is related to exactly two elements on the right and vice versa. However, the reductions yield games of a more restricted kind and this will be useful in Section 5.3. Say that a binary relation $R\subseteq A\times B$ is $2\leftrightarrow 2$ if it is the disjoint union of bipartite graphs $K_{2,2}$. That is to say $A$ and $B$ can be each partitioned into sets $A=\bigcup_{i}A_{i}$ and $B=\bigcup_{i}B_{i}$ so that each $A_{i}$ and $B_{i}$ has exactly two elements and $R=\bigcup_{i}A_{i}\times B_{i}$.
@@
Notice, that any $2\leftrightarrow 2$ constraint $\Phi$ can be represented by a pair of permutations $\pi_{1},\pi_{2}$, so that $\phi=\{(\pi_{1}(2\kappa+\iota_{1}),\pi_{2}(2\kappa+\iota_{2}))\mid 0\leq\kappa<\frac{q}{2},\iota_{1},\iota_{2}\in\{1,2\}\}$. We call such a constraint $D_{\pi_{1},\pi_{2}}$. Notice that any $2\leftrightarrow 2$ game must have an even alphabet. Also notice that multiple pairs of permutations can describe the same constraint.
@@ 4.1 Perfect completeness
Assume $I$ is a 3XOR instance on a set of variables $X$ that is satisfiable, and let $s:X\rightarrow\mathbb{F}_{2}$ be an assignment of values to the variables that satisfies it. Let $I^{w}_{2:2}$ denote the weighted $2$-to-$2$ game that $I$ maps to under the reduction. Then, for each vertex $(U,L)$ of $I^{w}_{2:2}$ the restriction of $s$ to $L+H_{U}$ is a valid label since all equations are satisfied, and it is easily seen that this labelling satisfies all constraints.
@@ 4.2 Vocabularies
For each positive integer $q$, we define a vocabulary $\tau_{\text{(T) 2-to-2}_{q}}$ such that structures in this vocabulary represent instances of transitive $2$-to-$2$ games over a label alphabet of size $q$. Let $S_{q}$ denote the collection of permutations of $[q]=\{1,\ldots,q\}$. Note that there is a natural bijective correspondence between $S_{q}$ and the $1$-to-$1$ relations on $[q]$. Now, let $S^{\#2}_{q}$ denote the set of pairs of permutations $(\pi_{1},\pi_{2})\in S_{q}\times S_{q}$ such that for all $i\in[q]$, $\pi_{1}(i)\neq\pi_{2}(i)$. Then, it is easily seen that each $2$-to-$2$ relation on $[q]$ can be seen (not uniquely) as the union of such a pair of permutations. Our vocabulary $\tau_{\text{(T) 2-to-2}_{q}}$ contains a binary relation for each element of $S_{q}$ and one for each element of $S^{\#2}_{q}$:
@@
For general (that is not necessarily edge distinct) $2$-to-$2$ games, we adopt a representation in which both vertices and edges are first-class elements. This allows for multiple edges between a pair of vertices, which may or may not all be associated with the same constraint. Specifically,
@@
where $C$ is unary, and the relations $\Phi_{\pi_{1},\pi_{2}}$ are all ternary. In a $\tau_{\text{(w) 2-to-2}_{q}}$-structure $\mathbb{A}$, the universe of $\mathbb{A}$ is the disjoint union of the set $V$ of vertices of $I^{w}_{2:2}$, and the set $C$ of constraints, with the unary relation $C$ picking out this set. For each $(\pi_{1},\pi_{2})\in S_{q}^{\#2}$, the relation $\Phi_{\pi_{1},\pi_{2}}\subseteq V^{2}\times C$ contains those triples $(u,v,c)$ where $c$ is a constraint linking $u$ and $v$ and $\Phi(c)$ is the $2$-to-$2$ relation associated with the pair $(\pi_{1},\pi_{2})$. We assume our structures satisfy the (first-order) axiom that ensures that for each $c$, there is exactly one pair $(u,v)$ and one relation $\Phi_{\pi_{1},\pi_{2}}$ in which the triple $(u,v,c)$ appears. Note that a structure in this vocabulary can also be understood as a edge distinct, wieghted game with integer weights. The weight of a constraint on an edge $(u,v)$ labelled with relation $R$ is simply the number of constraints $c$ for which $(u,v,c)$ is in $R$.
@@ 4.3 Undefinability of Regular 3XOR
The reduction in Section 3 starts from regular instances of 3XOR. In contrast, the undefinability result in Theorem 2.5 is stated for general 3XOR. Thus, we begin by arguing that the proof of Theorem 2.5 can actually be used to show the $\mathrm{FPC}$ undefinability of $\mathrm{Gap}\textsc{Regular3XOR}(1,\eta)$ for some $\eta$ strictly smaller than $1$.
We first note that $\mathrm{Gap}\textsc{3XOR}(1,\frac{1}{2}+\delta)$ is $\mathrm{FPC}$ undefinable even for “half-regular” 3XOR instances. That is, 3XOR instances where each variable appears in the same number of equations. To see this, note that [6, Lemma 5] uses a bipartite unique-neighbour expander graph with $r|X|$ nodes on the left and $|X|$ nodes on the right. Thus the graph is $3$-left-regular and is an ($\alpha|X|,\beta)$ expander. Such graphs exist for every $X$ by [29, Chapter 4]. By a variation of [29, Theorem 4.4], we claim the existence of such a graph with the extra condition that the graph is right-regular. Using this extra assumption on the graph in [6, Lemma 5] the proof establishes that $\mathrm{Gap}\textsc{3XOR}(1,\frac{1}{2}+\delta)$ is $\mathrm{FPC}$ undefinable even for “half-regular” 3XOR instances.
@@
A half-regular instance can be converted into a regular one by ensuring that any two equations share at most one variable.
@@
First, by the unique-neighbour expander property of the graph in [6, Lemma 5], we can assume that the half-regular 3XOR instance has no repeated equations or repeated variables within an equation. This half-regular instance $(X,\mathrm{Eq})$ can be converted into a regular one (call it $(X^{*},\mathrm{Eq}^{*})$) by replacing every equation $e:x+y+z=b$ with three equations (as done in [24]): $x+y_{e}+z_{e}=b,\quad x_{e}+y+z_{e}=b,\quad x_{e}+y_{e}+z=b$, where $x_{e},y_{e}$, and $z_{e}$ are new variables only used for these equations.
@@
$\pi^{U}(x,a_{1},a_{2},a_{3})\equiv(x=a_{1}\vee x=a_{2}\vee x=a_{3})\wedge(\mathrm{Eq}_{0}(a_{1},a_{2},a_{3})\vee\mathrm{Eq}_{1}(a_{1},a_{2},a_{3}))$
@@
{[$\wedge\{\left[\neg(x=x_{1}=x_{2}=x_{3})\wedge\neg(y=y_{1}=y_{2}=y_{3})\wedge(z=z_{1}=z_{2}=z_{3})\right]$
[$\vee\left[\neg(x=x_{1}=x_{2}=x_{3})\wedge(y=y_{1}=y_{2}=y_{3})\wedge\neg(z=z_{1}=z_{2}=z_{3})\right]$

### v2

@@ 3.3 The final (weighted) 2-to-2 game
The final step of the reduction is to transform the transitive game constructed in Section 3.2 into a weighted 22-to-22 game, getting rid of the 11-to-11 constraints. This weighted game is defined as follows.
Recall the transitive 22-to-22 game $\Theta(I)$ constructed in Section 3.2. The transitivity condition guarantees that the vertices of $\Theta(I)$ can be partitioned into cliques $C_{1},\dots,C_{m}$ so that edges in each clique are associated with 11-to-11 constraints. Moreover, these constraints are consistent in the sense that any colouring of a vertex $V$ in a clique $C$ can be extended in a unique way to a colouring of all vertices in $C$ so that all edge constraints in $C$ are satisfied. Also, by the transitivity condition, for distinct cliques $C_{i}$ and $C_{j}$, either all pairs $(u,v)\in C_{i}\times C_{j}$ are connected by 22-to-22 constraints or none are. Furthermore, these 22-to-22 constraints are consistent in the sense that given a clique-consistent colouring for $C_{i}$ and $C_{j}$, either all or none of these 22-to-22 constraints are satisfied.
@@
  • Choose a random pair $L,L^{\prime}$ so that $(U,L)$ and $(U,L^{\prime})$ are connected by a 22-to-22 edge. Let $C_{i}$ be the clique containing $(U,L)$ and $C_{j}$ be the clique containing $(U,L^{\prime})$
@@ 3.4 Irregular soundness case
For the result in Section 5.3, we need the $\mathrm{FPC}$-undefinability of a different gap problem based on 22-to-22 games. Specifically, we define the value of a game to be, not the fraction of constraints that can be satisfied, but the fraction of the vertices formed by the largest set $X$ so that all constraints between nodes in $X$ are satisfied. Moreover, we relax the notion of colouring to allow vertices to be coloured by multiple colours.
@@
For a 22-to-22 game $((V,E),\Sigma,\Phi)$, a colouring $c:V\rightarrow\binom{\Sigma}{j}$ satisfies a set $X\subseteq V$ if $\forall e\in E$, such that the vertices of $e$ are in $X$, $\exists a\in c(u),b\in c(v).(a,b)\in\Phi(e)$.
@@
For constants $j$ and $q$ define the function $\textsc{Irreg2to2}_{j,q}$ to take a 22-to-22 game $((V,E),\Sigma,\Phi)$ to the fraction $|X|/|V|$ where $X$ is the largest subset of $V$ that is satisfied by some $j$-colouring $c:V\rightarrow\binom{\Sigma}{j}$ .
@@
For every $\delta$ with $0<\delta<1$ and $j\in\mathbb{N}^{+}$, there exists $q\in\mathbb{N}^{+}$ so that $\mathrm{Gap}\textsc{Irreg2to2}_{j,q}(1,\delta)$ is not $\mathrm{FPC}$ definable.
Note that the value $\textsc{Irreg2to2}_{j,q}((V,E),\Sigma,\Phi)$ is unchanged if we simplify the game (i.e. remove edge weights and multiplicities of the same constraint between a pair of nodes), therefore Theorem 3.5 applies even in restriction to simple, unweighted 2-to-2 games.
@@
For a weighted simple 22-to-22 game $I=((V,E),\Sigma,\Phi)$ as constructed in Section 3.3 with $q=|\Sigma|$, if $\textsc{Irreg2to2}_{j,q}((V,E),\Sigma,\Phi)\geq\delta$, then $\textsc{Weight2to2}(I)\geq\frac{\gamma(\delta)}{j^{2}}$ for some $\gamma(\delta)\in\Omega(\delta^{2})$.
@@
Let $c$ be a $j$-colouring of $V$ that satisfies a set $X$ with $|X|/|V|\geq\delta$. By [26, Remark 3.4.9], there is a $\gamma(\delta)\in\Omega(\delta^{2})$ (weighted) fraction of the edges $E$ which are satisfied by $c$, in the sense that for each such edge $e$ on vertices $u,v$ there are colours $a$ and $b$ in $c(u)$ and $c(v)$ respectively such that $(a,b)\in\Phi(e)$. We now construct a standard colouring by a random process. That is, for each vertex $v\in V$, independently choose a colour $\chi(v)$ from $c(v)$ uniformly at random. For an edge $e$ on vertices $u,v$, let $\Xi(e)$ be the indicator variable indicating whether $(\chi(u),\chi(v))\in\Phi(e)$ and let $\Xi$ be the overall value of the colouring $\chi$. If $(u,v)\in X^{2}$, the probability that $\chi$ satisfies the constraint $\Phi(e)$ is at least $\frac{1}{j^{2}}$, as by definition, among the $j^{2}$ pairs in $c(u)\times c(v)$, at least one satisfies the constraint. Then, we have,
@@
$\geq\sum_{e\in E|_{X}}w(e)\mathbb{E}[\Xi(e)]\geq\sum_{e\in E|_{X}}w(e)\frac{1}{j^{2}}\geq\gamma(\delta)\frac{1}{j^{2}}$
where $E|_{X}$ denotes the subset of $E$ containing those edges both of whose vertices are in $X$. Thus, there is a colouring that satisfies at least an $\Omega(\delta^{2})$ (weighted) fraction of the constraints. ∎
From Lemma 3.6, we can conclude Theorem 3.5. Indeed, fix a value of $\delta$ and $j$ and let $\gamma=\gamma(\delta)$ be the value determined by Lemma 3.6. Then, the proof of Theorem 3.1 gives us a $q$, and $\eta$ and an $\mathrm{FO}$ reduction that maps satisfiable 3XOR instances to satisfiable 22-to-22 games, and maps instances that are at most $\eta$-satisfiable to 22-to-22 games with value less than $\frac{\gamma}{j^{2}}$. Then, by Lemma 3.6, this same reduction also maps at most $\eta$-satisfiable 3XOR instances to 2-to-2 games $I$ for which $\textsc{Irreg2to2}_{j,q}(I)<\delta$.
@@ 3.5 $2\leftrightarrow 2$ games
The definition of 22-to-22 games, Definition 2.4 only requires each constraint $\Phi(u,.v)$ to be a 22-to-22 relation, meaning that each element on the left is related to exactly two elements on the right and vice versa. However, the reductions yield games of a more restricted kind and this will be useful in Section 5.3. Say that a binary relation $R\subseteq A\times B$ is $2\leftrightarrow 2$ if it is the disjoint union of bipartite graphs $K_{2,2}$. That is to say $A$ and $B$ can be each partitioned into sets $A=\bigcup_{i}A_{i}$ and $B=\bigcup_{i}B_{i}$ so that each $A_{i}$ and $B_{i}$ has exactly two elements and $R=\bigcup_{i}A_{i}\times B_{i}$.
@@
Notice, that any $2\leftrightarrow 2$ constraint $\Phi$ can be represented by a pair of permutations $\pi_{1},\pi_{2}$, so that $\phi=\{(\pi_{1}(2\kappa+\iota_{1}),\pi_{2}(2\kappa+\iota_{2}))\mid 0\leq\kappa<\frac{q}{2},\iota_{1},\iota_{2}\in\{1,2\}\}$ (for a fixed $\kappa$, $\pi_{1}(2\kappa),\pi_{1}(2\kappa+1),\pi_{2}(2\kappa),\pi_{2}(2\kappa+1)$ form a $K_{2,2}$ component in the constraint graph). We call such a constraint $D_{\pi_{1},\pi_{2}}$. Notice that any $2\leftrightarrow 2$ game must have an even alphabet. Also notice that multiple pairs of permutations can describe the same constraint.
@@ 4.1 Perfect completeness
Assume $I$ is a 3XOR instance on a set of variables $X$ that is satisfiable, and let $s:X\rightarrow\mathbb{F}_{2}$ be an assignment of values to the variables that satisfies it. Let $I^{w}_{2:2}$ denote the weighted 22-to-22 game that $I$ maps to under the reduction. Then, for each vertex $(U,L)$ of $I^{w}_{2:2}$ the restriction of $s$ to $L+H_{U}$ is a valid label since all equations are satisfied, and it is easily seen that this labelling satisfies all constraints.
@@ 4.2 Vocabularies
For each positive integer $q$, we define a vocabulary $\tau_{\text{(T) 2-to-2}_{q}}$ such that structures in this vocabulary represent instances of transitive 22-to-22 games over a label alphabet of size $q$. Let $S_{q}$ denote the collection of permutations of $q]=\{1,\ldots,q\}$. Note that there is a natural bijective correspondence between $S_{q}$ and the 11-to-11 relations on $[q]$. Now, let $S^{\#2}_{q}$ denote the set of pairs of permutations $(\pi_{1},\pi_{2})\in S_{q}\times S_{q}$ such that for all $i\in[q]$, $\pi_{1}(i)\neq\pi_{2}(i)$. Then, it is easily seen that each 22-to-22 relation on $[q]$ can be seen (not uniquely) as the union of such a pair of permutations. In particular, each $i$ is related to exactly the two elements $\pi_{1}(i)$ and $\pi_{2}(i)$. Our vocabulary $\tau_{\text{(T) 2-to-2}_{q}}$ contains a binary relation for each element of $S_{q}$ and one for each element of $S^{\#2}_{q}$:
@@
For general (that is not necessarily edge distinct) 22-to-22 games, we adopt a representation in which both vertices and edges are first-class elements. This allows for multiple edges between a pair of vertices, which may or may not all be associated with the same constraint. Specifically,
@@
where $C$ is unary, and the relations $\Phi_{\pi_{1},\pi_{2}}$ are all ternary. In a $\tau_{\text{(w) 2-to-2}_{q}}$-structure $\mathbb{A}$, the universe of $\mathbb{A}$ is the disjoint union of the set $V$ of vertices of $I^{w}_{2:2}$, and the set $C$ of constraints, with the unary relation $C$ picking out this set. For each $(\pi_{1},\pi_{2})\in S_{q}^{\#2}$, the relation $\Phi_{\pi_{1},\pi_{2}}\subseteq V^{2}\times C$ contains those triples $(u,v,c)$ where $c$ is a constraint linking $u$ and $v$ and $\Phi(c)$ is the 22-to-22 relation associated with the pair $(\pi_{1},\pi_{2})$. We assume our structures satisfy the (first-order) axiom that ensures that for each $c$, there is exactly one pair $(u,v)$ and one relation $\Phi_{\pi_{1},\pi_{2}}$ in which the triple $(u,v,c)$ appears. Note that a structure in this vocabulary can also be understood as a edge distinct, wieghted game with integer weights. The weight of a constraint on an edge $(u,v)$ labelled with relation $R$ is simply the number of constraints $c$ for which $(u,v,c)$ is in $R$.
@@ 4.3 Undefinability of Regular 3XOR
The reduction in Section 3 starts from regular instances of 3XOR. In contrast, the undefinability result in Theorem 2.5 is stated for general 3XOR. Thus, we begin by arguing that the proof of Theorem 2.5 can actually be used to show the $\mathrm{FPC}$ undefinability of $\mathrm{Gap}\textsc{Regular3XOR}(1,\eta)$ for some $\eta$ strictly smaller than $1$. We break down the proof into two parts.
Step 1: Undefinability of half-regular 3XOR
Definition 4.1.
@@
A 3XOR instance is $d$-half-regular if every variable appears in at most $d$ equations.
Lemma 4.2.
$\mathrm{GapDHalfRegular3XOR}(1,\frac{1}{2}+\delta)$ is $\mathrm{FPC}$-undefinable for any $\delta>0$, $d>0$
In order to see this, we can modify the proof of Theorem 2.5 as given in [7] to apply to half-regular 3XOR instances. In order to show this, we need to quickly sketch the proof from [7] to see how to alter it to get the strenghtened result:
Theorem 2.5 is obtained as a direct consequence of the following four lemmas. The statements are quoted verbatim from [7] and we provide some explanation of the terminology below.
Lemma 4.3 ([7], Lemma 2).
For every 3XOR instance $I$ and every integer $k\geq 3$, if $I$ is $k$-locally satisfiable, then
$G(I)\equiv_{C_{k}}G(I_{0}).$
Lemma 4.4 ([7], Lemma 3).
For every 3XOR instance $I$ and every $c,s\in[0,1]$, the following hold:
  1. 1.
    If $I$ is $c$-satisfiable, then $G(I)$ is $c$-satisfiable.
  2. 2.
    If $I$ is not $s$-satisfiable, then $G(I)$ is not $\left(\frac{1}{2}+\frac{s}{2}\right)$-satisfiable.
Lemma 4.5 ([7], Lemma 4).
For every two reals $\varepsilon>0$ and $\delta>0$ there exists an integer $c>0$ such that for every sufficiently large integer $n$ and every matrix
$A\in\{0,1\}^{m\times n},$
where $m=cn$ and each row of $A$ has exactly three ones, if $b$ is chosen uniformly at random in $\{0,1\}^{m}$, then, with probability at least $1-\delta$, both 3XOR instances $Ax=b$ and $G(Ax=b)$ are at most $\left(\frac{1}{2}+\varepsilon\right)$-satisfiable.
Note that a 3XOR instance can be expressed as the equation $Ax=b$, where $A$ is a matrix with every row representing an equation, every column representing a variable, and $A[i,j]=1$ if, and only if variable $j$ appears in equation $i$, while $b$ is the vector of equation right-hand sides.
Lemma 4.6 ([7], Lemma 5).
For every integer $r>0$ there is a real $\gamma>0$ such that for every sufficiently large integer $n$ there is a matrix
$A\in\{0,1\}^{m\times n},$
where $m=rn$, such that each row of $A$ has exactly three ones and, for every vector $b\in\{0,1\}^{m}$, the 3XOR instance $Ax=b$ is $k$-locally satisfiable for
$k\leq\gamma n.$
In order to understand these lemmas we need to define some terminology. First, we define the map $G$ referred to in Lemmas 4.3 and 4.4. It takes a 3XOR instance $I$ and replaces every equation $x+y+z=b$ with 8 equations of the form $x_{i}+y_{j}+z_{k}=b+i+j+k$ for all possible triplets $(i,j,k)\in\{0,1\}^{3}$. By construction, it is easy to see that $G$ maps any half-regular 3XOR instance to another half-regular instance (any $x_{i}$ appears in $4$ times as many equations in $G(I)$ as the number of equations $x$ appears in $I$). Given a 3XOR instance $I:Ax=b$, $I_{0}$ is defined as $Ax=0$. We omit a rigorous definition fo the notion of $k$-local satisfiability referred to in Lemmas 4.3 and 4.6 as it is not needed. We just note that the $k$-locally satisfiable of the instance $Ax=b$ in Lemma 4.6 is established directly from the expansion properties of the bipartite graph for which the $A$ is the biadjacency matrix. We return to these expansion properties below.
Lemmas 4.34.4 and 4.5 hold for all 3XOR instances, including half-regular ones. So in order to establish Lemma 4.2, it suffices to show that in the proof of Lemma 4.6, the matrix $A$ can be constructed to satisfy the additional condition that every column has the same number of $1$s (i.e. each variable appears in $3r$ equations). Given this, Lemma 4.2 follows, since we then have a family of instances of the form $G(Ax=b)$ which are no more than $(\frac{1}{2}+\epsilon)$-satisfiable and a family of instances of the form $G(Ax=0)$ which are satisfiable. These families are half-regular by construction and not separable by any class of bounded counting width by Lemma 4.3.
We mentioned above that the $k$-locally satisfiable of the instance $Ax=b$ is established by taking $A$ to be the biadjacency matrix of an arbitrary bipartite unique-neighbour expander graph with parameters $(rn,n,3,\alpha n,\beta)$. This means that it is a bipartite graph with $rn$ nodes on the left, $n$ nodes on the right, which is $3$-left-regular, and an $(\alpha n,\beta)$ unique neighbour expander, which we define next.
Definition 4.7.
We call a bipartite graph an $(X,Y)$ unique-neighbour expander if for all sets of left-nodes $S$ with size at most $X$, the set of unique neighbours, $N(S)\triangleq\{v\mid\exists!u\in S.(u,v)\in E\}$ has size at least $Y|S|$.
Thus, to complete the proof, we just need to show that there are arbitrarily large such graphs which are also right regular. Specifically, we need to show that for some fixed $r\in\mathbb{N^{+}},\alpha,\beta>0$, there exists a bipartite unique neighbour expander that
  • has $rn$ left nodes and $n$ right nodes;
  • is $3$-left-regular and $3r$-right-regular; and
  • is a unique-neighbour $(\alpha n,\beta)$-expander.
We show this (using $\beta=1$) by a slight variation of the proof of [30, Theorem 4.4] (case $D=3$):
Suppose we generate a 3-left-regular, $3r$-right-regular bipartite graph with $rn$ left nodes and $n$ right nodes with the following random process
  1. 1.
    Start with an empty bipartite graph with $rn$ left nodes and $n$ right nodes. Randomly partition the left nodes into $r$ groups of $n$.
  2. 2.
    In $3r$ rounds, do the following: take one of the $r$ groups on the left (taking each group exactly $3$ times) and generate a random matching between the group and the nodes on the right, which has no common edge with the previously generated matchings.
  3. 3.
    The edges of our resulting graph are the union of these $3r$ matchings.
Lemma 4.8.
A graph generated with the random process above is a $(\alpha n,1)$ unique-neighbour expander with probability at least $\frac{1}{2}$ for some $\alpha$.
Proof.
Fix a subset $S$ of the left vertices of size $K\leq\alpha n$. List the $3K$ neighbours of $S$ (including multiplicites, i.e. if a right node $v$ has $l$ incoming edges from $S$, it is listed $l$ times) in an arbitrary order, naming them $V_{1},V_{2},...,V_{3K}$. Also let $U_{1},U_{2},...U_{3K}$ be elements of $S$, each element mentioned three times, such that there is an edge $(U_{i},V_{i})$ for all $i$. The probability of $V_{i}$ not being a unique neighbour is at most $\sum_{i\neq j}P(V_{i}=V_{j})$. Given $U_{i}$, $U_{j}$ and the round in which the corresponding edges have been selected in the random process, we know that for all $i\neq j$ $P(V_{i}=V_{j})$ is
  • 0, if $U_{i}=U_{j}$, or the nodes $U_{i}$ and $U_{j}$ belong to the same group of $n$, and the edges $(U_{i},V_{i})$ and $(U_{j},V_{j})$ are selected in the same round (recall that we select matchings in a round, so the edges must be disjoint)
  • $\frac{1}{n}$, if $U_{i}$ and $U_{j}$ belong to different groups of $n$ (the edge choices are independent)
  • $\frac{1}{n-1},$ if $U_{i}$ and $U_{j}$ belong to the same group but the edges are chosen in a different round. To see this, say the edge $(U_{i},V_{i})$ is selected in round $k$. Then in the same round, an edge $(U_{l},V_{l})$ is selected such that $U_{l}=U_{j}$. Since in every round, we choose a matching, $(V_{l}\neq V_{i})$, $V_{j}$ must be different from $V_{l}$, and with no further constraint, it can be any of the remaining n1n-1 right-nodes, with uniform probability.
Thus,
$P(V_{i}\text{ is not a unique neighbour})\leq\sum_{i\neq j}P(V_{i}=V_{j})\leq\frac{3K-1}{N-1}\leq\frac{3K}{N}$
Note that the rest of the calculation is identical to [30, Theorem 4.4], and the explanation from here is an almost verbatim reproduction of the proof from the book, substituting $D=3$:
  • The probability that $S$ violates the expander property is
[…]
    $\leq\binom{3K}{2K}\left(\frac{3K}{N}\right)^{2K}.$
  • Let $p_{K}$ be the probability that any set of size $K$ does not expand with $\beta=1$. Then
    $p_{K}\leq\binom{N}{K}\binom{KD}{2K}\left(\frac{KD}{N}\right)^{2K}$
    $\leq\left(\frac{Ne}{K}\right)^{K}\left(\frac{KDe}{2K}\right)^{2K}\left(\frac{KD}{N}\right)^{2K}=\left(\frac{e^{3}D^{4}K}{4N}\right)^{K}.$
  • Noting $K\leq\alpha N$, let $\alpha=\frac{1}{e^{3}D^{4}}$, so $p_{K}\leq 4^{-K}$. Thus,
    $P\bigl[G\text{ is not a unique-neighbour }(\alpha N,1)\text{ expander}\bigr]\leq\sum_{K=1}^{\lfloor\alpha N\rfloor}4^{-K}<\frac{1}{2}.$
This shows that such expander graph exists, completing the proof of the undefinability gap for half–regular 3XOR.
Step 2: Undefinability of regular 3XOR
A half-regular instance can be converted via a reduction into a regular one by ensuring that any two equations share at most one variable.
@@
A half-regular instance $(X,\mathrm{Eq})$ can be converted into a regular one (call it $(X^{*},\mathrm{Eq}^{*})$) by replacing every equation $e:x+y+z=b$ with three equations (as done in [26]): $x+y_{e}+z_{e}=b,\quad x_{e}+y+z_{e}=b,\quad x_{e}+y_{e}+z=b$, where $x_{e},y_{e}$, and $z_{e}$ are new variables only used for these equations.
@@
$\pi^{U}(x,a_{1},a_{2},a_{3})\equiv[(x=a_{1}\vee x=a_{2}\vee x=a_{3})\wedge(\mathrm{Eq}_{0}(a_{1},a_{2},a_{3})\vee\mathrm{Eq}_{1}(a_{1},a_{2},a_{3}))]\vee(x=x_{1}=x_{2}=x_{3})$
@@
$\wedge\{\left[\neg(x=x_{1}=x_{2}=x_{3})\wedge\neg(y=y_{1}=y_{2}=y_{3})\wedge(z=z_{1}=z_{2}=z_{3})\right]$
$\vee\left[\neg(x=x_{1}=x_{2}=x_{3})\wedge(y=y_{1}=y_{2}=y_{3})\wedge\neg(z=z_{1}=z_{2}=z_{3})\right]$
@@
Explanation: we define any variables $x_{e}$ using $x$ and the three variables of the equations. We represent original variables $x$ by four copies of $x$,